# On the existence of unbiased resilient estimators in discrete quantum systems

### v3

@@ I Introduction
Recent efforts have aimed to derive tighter bounds for quantum parameter estimation. For example, Gessner et al. [30] proposed a hierarchy of lower bounds on estimator variance by imposing additional unbiasedness constraints. This hierarchy includes the quantum CRB (QCRB) [31] as its lowest-order bound, while higher-order cases include quantum analogs of the Barankin, Bhattacharyya, and Abel bounds.
In this work, we analyze both classical and quantum Bhattacharyya bounds (QBhBB) and assess their performance relative to the CRB and Bayesian approaches, particularly in scenarios with uncertain prior knowledge of the parameter. We first illustrate these bounds with two paradigmatic cases, namely, the estimation of the variance of a Gaussian distribution, and the analysis of the performance of a Mach-Zehnder interferometer setup. Our findings reveal that the BhBB-inspired strategy can provide a significant advantage in specific regimes. Furthermore, we investigate the conditions under which the classical Bhattacharyya bound becomes non-computable, deriving the necessary and sufficient criteria for the existence of these bounds. Notably, we demonstrate that the quantum Bhattacharyya bound has more relaxed conditions for computability compared to its classical counterpart.
@@ II Estimation theory
To extract information from a physical system, measurements must be performed. The measurement outcome is modeled as a random variable that follows a probability distribution dependent on the state of the system. The information is encoded in this random variable, and estimation theory is the branch of statistics that addresses optimal protocols and the ultimate limits for information extraction. The Cramér-Rao bound (CRB) establishes a limit on the precision of estimating a quantity or parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distributions $P_{\theta}(x)$, parameterized by the parameter $\theta$, which is being estimated. An estimator $\tilde{\Theta}(x)$ is a function of the random measurement outcome that provides an estimate of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE), and variance $(\Delta\tilde{\Theta})^{2}$.
@@
The mean value is defined as $\langle f\rangle_{\theta}=\sum_{x}f(x)P_{\theta}(x)\>$. An estimator $\tilde{\Theta}_{\theta_{0}}(x)$ is locally unbiased on the point $\theta_{0}$ when $\langle\tilde{\Theta}_{\theta_{0}}\rangle_{\theta_{0}}=\theta_{0}$ and $\frac{db(\theta)}{d\theta}|_{\theta=\theta_{0}}=0$. Note that for unbiased estimators MSE and the variance coincide. Unbiased Cramér-Rao, usually named as the Cramér-Rao bound (CRB), states that for any unbiased estimator
@@ II.1 Bhattacharyya bound
Bhattacharyya bounds (BhB) were introduced in the context of classical frequentist parameter estimation problems to provide a lower bound for the variance of an estimate [26]. The BhB gives a tighter bound than the Cramér-Rao bound (CRB) by incorporating higher-order derivatives of the probability distribution in the calculation. It is shown that the BhB converges to the variance of the best unbiased estimator—the estimator that is unbiased for any point in the region of the allowed values of the parameter—when the sampling distribution belongs to an exponential family of distributions [34].
@@
and $\bm{g}(x)^{T}=(\partial P_{\theta_{0}}(x),\partial^{2}P_{\theta_{0}}(x),...,\partial^{n}P_{\theta_{0}}(x))$. We would like to point out the similarities between the $n^{th}$order BhB bound and the multiparameter Cramér-Rao bound (MCRB) [35] in a multiparameter estimation problem with $n$ parameters. The BhB can be codified as a particular case of the CRB in a multiparameter estimation setting.
@@ II.2 Application: Gaussian Distribution
We consider the problem of estimating the variance $\sigma$ of a zero mean Gaussian probability distribution,
@@
To address this, we adopt Jeffreys prior, which is given by $p_{\text{prior}}(\theta)\propto\sqrt{F_{C}}$, where $F_{C}$ is the classical Fisher information. Jeffreys prior is particularly useful as it remains invariant under reparameterization, ensuring consistency across different representations of the parameter. In Fig. [1], we depict the mean squared error (MSE) and the bias of the different estimators. As expected, the Bayesian approach gives a lower MSE over a large range of values of $\sigma$, but its bias is the largest. On the other hand, the estimator satisfying the order-5 Bhattacharyya bound (BhB) has the smallest bias in the interval considered, but its MSE increases rapidly for large values of $\sigma$. The most suitable estimator is the one saturating the order-2 BhB, as its MSE remains more constant than $\text{MSE}_{\text{CR}}$, with a bias close to zero.
@@ II.3 Quantum bounds
From Eq. (13) it is deduced that the SLD is a hermitian operator and its eigenvectors form an orthonormal basis. In the derivation of the QCRB, a general POVM $\{\Pi_{x}\}$ was considered but it was shown that a projection on the basis of eigenvectors of $L$ is the optimal measurement that saturates the QCRB [14]. As in the classical estimation theory, the QCRB is always saturable in the asymptotic regime. Furthermore, it has recently been shown to be non-saturable within a Bayesian perspective [37, 24]. The BhB can be extended to quantum estimation theory [30] in a coherent way that includes them as a specific case of a larger set of bounds. Doing so, the QBhB state that the variance of any unbiased estimator with null $n$ first bias derivatives satisfies
@@
Again, the last equality holds when $\bm{Q}^{-1}$ exists. The $n\times n$ matrix $\bm{Q}$ is defined as
@@
$\bm{\lambda}^{\top}\bm{Q}^{-1}\bm{g}(x)=P_{\theta_{0}}(x)\left(\tilde{\Theta}_{\theta_{0}}(x)-\theta_{0}\right).$ (17)
@@ III.1 Classical
where we defined $\partial^{l}P_{\theta_{0}}(x)\equiv\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}\big{|}_{\theta=\theta_{0}}$. The system has $n+1$ equations and $N$ variables. We consider the case where the first $N-1$ derivatives are linearly independent. This implies that if $n\leq N-1$, there is always a solution.
@@
It is then concluded that for a probability distribution with $N$ points, it is advisable to compute the BhB bound up to order $N-1$. In the appendix (A), the conditions for the existence of the estimator are given. It is also shown that the bound diverges if and only if the estimator does not exist.
@@ III.2 Quantum
We extend the previous results to determine when computing QBhB gives more information about the estimation problem. In quantum estimation theory the measurement is a POVM $\{\Pi_{i}\}$, where $i$ are the possible outcomes, each one with probability $P_{\theta}(i)=\operatorname{Tr}(\Pi_{i}\rho_{\theta})$. The classical processing of the measurement outcomes is accounted by the estimator $\tilde{\Theta}(x)$ that associates an estimate $\tilde{\Theta}(i)$ to the outcome $i$. Every estimate $\tilde{\Theta}(i)$ has a probability given by $P_{\theta}(i)$. Both tasks, measurement, and classical processing, can be treated simultaneously as a hermitian operator $\hat{\Theta}=\sum_{i}\tilde{\Theta}(i)\Pi_{i}$ [23]. The conditions of Eq. (LABEL:condiciones) are now written in terms of the operator $\hat{\Theta}$ as
@@
$\vec{x}=$ $\begin{pmatrix}\hat{\Theta}_{1}\\ \hat{\Theta}_{2}\\ \vdots\\ \hat{\Theta}_{\frac{N(N+1))}{2}}\end{pmatrix},\quad\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix},$
So we have a system of equations equivalent to the one for a classical probability distribution with $\frac{N(N+1)}{2}$ support points. As before if this system of equations cannot be solved the QBhB variance will go to infinity. On the other hand, if the first $\frac{N(N+1)}{2}-1$ derivatives are linearly independent, for all $n>\frac{N(N+1)}{2}-1$ one has that $n^{th}$order QBhB is equal to $\frac{N(N+1)}{2}-1$ order QBhB. The difference between the quantum and classical cases is that in the former we have the freedom to choose a measurement. This extra freedom implies that the highest relevant QBhB bound –which gives a different bound than lower orders– is lifted in the quantum case because $N-1<\frac{N(N+1)}{2}-1$. In appendix (B) an example is studied in detail.
@@
To resolve this apparent contradiction, we explain why the $N$-order quantum Bhattacharyya bound (QBhB) need not be equal to the $(N-1)$-order QBhB. The key point is that, to compute the $N$- and $(N+1)$-order QBhB, different measurements have to be performed, resulting in different probability distributions $P_{N}$ and $P_{N+1}$. This extra freedom in the choice of measurements allows quantum systems to access more non-trivial information. In other words, the maximal order for which the QBhB provides distinct information is higher than in classical scenarios. Thus, since $P_{N}$ and $P_{N+1}$ are different probability distributions, there is no discrepancy with what we presented in Section (III.1).
@@ IV Mach-Zehnder interferometer
Historically, greater attention has been given to the Barankin bounds [27, 28]. In contrast, to the best of our knowledge, Bhattacharyya bounds have not been computed for any physical problem before. In this work, we compute the Bhattacharyya bounds and identify the estimators that achieve these bounds for the paradigmatic Mach-Zehnder interferometer. We demonstrate that, when the prior distribution is not sharply peaked and we move away from the point where unbiased estimators are constructed, the Bhattacharyya bounds provide an improvement in the achievable precision.
@@ IV.0.1 Mach-Zehnder
In the Mach-Zehnder configuration, depicted in Fig. (2), two modes of the electromagnetic field interfere on a balanced beamsplitter, the output beams acquire a relative phase $\theta=\varphi_{a}-\varphi_{b}$, and interfere again on the second beamsplitter. Finally, the photon numbers $n_{a}$ and $n_{b}$ are measured at the output ports.
@@
where $J_{q}(x)$ are the Bessel functions. We compute the CRB, the $2^{\text{nd}}$-order BhB, and the estimators achieving each bound at $\theta_{0}=10^{-3}$. In Fig. (3), we compare the mean squared error (MSE) of these two estimators within an interval of $\theta$ that contains $\theta_{0}$.
One can conclude that when the prior knowledge is not sharp, the BhB offers a more robust bound; i.e., $\text{MSE}_{\text{Bh}}$ remains flat over an interval of $\theta$, whereas $\text{MSE}_{\text{CR}}$ increases rapidly when $\theta\neq\theta_{0}$. In addition, the protocol that saturates the BhB achieves a smaller MSE. Therefore, it may be advisable to use this bound as a figure of merit instead of the CRB in certain specific scenarios.
(a) MSE
(b) Bias
@@
In Fig. (4), we analytically compute the second-order Bhattacharyya bound and the Cramér-Rao bound as a function of the increasing number of photons. We observe that the Cramér-Rao bound achieves the Heisenberg limit, as expected. Interestingly, the Bhattacharyya bound also reaches the Heisenberg limit, indicating that a tighter bound can still achieve Heisenberg scaling. This result suggests an alternative approach to constructing estimators that can achieve optimal scaling even with limited prior knowledge of the true parameter value. In particular, this demonstrates that it is possible to design estimation protocols that are more robust to uncertainties in prior knowledge while still reaching the Heisenberg limit in precision.
Figure 4: A logarithmic plot of the Cramér-Rao bound and the second-order Bhattacharyya bound is shown. Both bounds scale as $\sim\frac{1}{n^{2}}$, which means that both bounds attain the Heisenberg limit.

### v4

# Existence of unbiased estimators in discrete quantum systems

@@ I Introduction
Recent efforts have aimed to derive tighter bounds for quantum parameter estimation. The first result on this line was published by Tsuda [30]. More recently, Gessner and Smerzi [31] proposed a hierarchy of lower bounds on estimator variance by imposing additional unbiasedness constraints. This hierarchy includes the quantum CRB (QCRB) [32] as its lowest-order bound, while higher-order cases include quantum analogs of the Barankin, Bhattacharyya, and Abel bounds.
In this work, we analyze both classical and quantum Bhattacharyya bounds (QBhBB) and assess their performance relative to the CRB and Bayesian approaches, particularly in scenarios with uncertain prior knowledge of the parameter. We first illustrate these bounds with two paradigmatic cases, namely, the estimation of the variance of a Gaussian distribution, and the analysis of the performance of a Mach-Zehnder interferometer setup. Our findings reveal that the BhB-inspired strategy can provide a significant advantage in specific regimes. Furthermore, we investigate the conditions under which the classical Bhattacharyya bound becomes non-computable, deriving the necessary and sufficient criteria for the existence of these bounds. Notably, we demonstrate that the quantum Bhattacharyya bound has more relaxed conditions for computability compared to its classical counterpart.
@@ II Estimation theory
To extract information from a physical system, measurements must be performed. The measurement outcome is modeled as a random variable that follows a probability distribution dependent on the state of the system. The information is encoded in this random variable, and estimation theory is the branch of statistics that addresses optimal protocols and the ultimate limits for information extraction. The Cramér-Rao bound (CRB) establishes a limit on the precision of estimating a quantity or parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distributions $P_{\theta}(x)$, parametrized by the parameter $\theta$, which is being estimated. An estimator $\tilde{\Theta}(x)$ is a function of the random measurement outcome that provides an estimate of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE), and variance $(\Delta\tilde{\Theta})^{2}$.
@@
The mean value is defined as $\langle f\rangle_{\theta}=\sum_{x}f(x)P_{\theta}(x)\>$. An estimator $\tilde{\Theta}_{\theta_{0}}(x)$ is locally unbiased on the point $\theta_{0}$ when $\langle\tilde{\Theta}_{\theta_{0}}\rangle_{\theta_{0}}=\theta_{0}$ and $\frac{db(\theta)}{d\theta}|_{\theta=\theta_{0}}=0$. The subindex $\theta_{0}$ is used to remind that the estimator is unbiased at that point. Note that for unbiased estimators the MSE and the variance coincide. Unbiased Cramér-Rao, usually named as the Cramér-Rao bound (CRB), states that for any unbiased estimator
@@ II.1 Bhattacharyya bound
Bhattacharyya bounds (BhB) were introduced in the context of classical frequentist parameter estimation problems to provide a lower bound for the variance of an estimate [26]. The BhB gives a tighter bound than the Cramér-Rao bound (CRB) by incorporating higher-order derivatives of the probability distribution in the calculation. It is shown that the BhB converges to the variance of the best unbiased estimator—the estimator that is unbiased for any point in the region of the allowed values of the parameter—when the probability distribution belongs to an exponential family of distributions [35].
@@
and $\bm{g}(x)^{T}=(\partial P_{\theta_{0}}(x),\partial^{2}P_{\theta_{0}}(x),...,\partial^{n}P_{\theta_{0}}(x))$.
@@ II.2 Application: Gaussian Distribution
We consider the problem of estimating the variance $\sigma^{2}$ of a zero mean Gaussian probability distribution,
@@
To address this, we adopt Jeffreys prior, which is given by $p_{\text{prior}}(\theta)\propto\sqrt{F_{C}}$, where $F_{C}$ is the classical Fisher information. Jeffreys prior is particularly useful as it remains invariant under reparameterization, ensuring consistency across different representations of the parameter. In Fig. 1, we depict the mean squared error (MSE) and the bias of the different estimators. As expected, the Bayesian approach gives a lower MSE over a large range of values of $\sigma$, but its bias is the largest. On the other hand, the estimator satisfying the order-5 Bhattacharyya bound (BhB) has the smallest bias in the interval considered, but its MSE increases rapidly for large values of $\sigma$.
@@ II.3 Quantum bounds
From Eq. (13) it is deduced that the SLD is a hermitian operator and its eigenvectors form an orthonormal basis. In the derivation of the QCRB, a general POVM $\{\Pi_{x}\}$ was considered but it was shown that a projection on the basis of eigenvectors of $L$ is the optimal measurement that saturates the QCRB [14]. As in the classical estimation theory, the QCRB is always saturable in the asymptotic regime. Furthermore, it has recently been shown to be non-saturable within a Bayesian perspective [37, 24]. BhB can be expanded to its quantum counterpart QBhB, as an example of bounds in quantum estimation theory [31]. Doing so, the QBhB state that the variance of any unbiased estimator with null $n$ first bias derivatives satisfies
@@
Again, the last equality holds when $\bm{Q}^{-1}$ exists. In Ref. [38] the authors study the consequences of non invertible matrices in quantum multiparameter metrology. The $n\times n$ matrix $\bm{Q}$ is defined as
@@
$\bm{\lambda}^{\top}\bm{Q}^{-1}\bm{g}(x)=P_{\theta_{0}}(x)\left(\tilde{\Theta}_{\theta_{0}}(x)-\theta_{0}\right),$ (17)
where as in the classical case $\bm{g}(x)^{T}=(\partial P_{\theta_{0}}(x),\partial^{2}P_{\theta_{0}}(x),...,\partial^{n}P_{\theta_{0}}(x))$.
@@ III.1 Classical
where we defined $\partial^{l}P_{\theta_{0}}(x)\equiv\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}\big{|}_{\theta=\theta_{0}}$. The system has $n+1$ equations and $N$ variables. Moreover, we define $A_{j}$ as the vector made of the elements of the $j$-th row of $A$, i.e.
$A_{j}=\begin{pmatrix}\partial^{j}P_{\theta_{0}}(x_{1})&\partial^{j}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{j}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}.$
We now consider the case where the set of vectors $\{A_{j}:1<j\leq N\}$ is linearly independent. This implies that if $n\leq N-1$, there is always a solution.
@@
It is then concluded that for a probability distribution with $N$ points, it is advisable to compute the BhB bound up to order $N-1$. In the appendix A, the conditions for the existence of the estimator are given. It is also shown that the bound diverges if and only if the estimator does not exist.
@@ III.2 Quantum
We extend the previous results to determine how the maximal degree –that gives nontrivial information– of the QBhB is limited by the dimension of quantum system. The consequences of finite-dimensional systems in parameter estimation have also been studied Ref. [39, 40]. In quantum estimation theory the measurement is a POVM $\{\Pi_{i}\}$, where $i$ are the possible outcomes, each one with probability $P_{\theta}(i)=\operatorname{Tr}(\Pi_{i}\rho_{\theta})$. The classical processing of the measurement outcomes is accounted by the estimator $\tilde{\Theta}(x)$ that associates an estimate $\tilde{\Theta}(i)$ to the outcome $i$. Every estimate $\tilde{\Theta}(i)$ has a probability given by $P_{\theta}(i)$. From now on we are going to consider projective value measurement (PVM). These measurements are optimal in the sense that they saturate the CRB, i.e. there is always a projective measurement that satisfies $F_{C}=F_{Q}$. This measurement consists in projecting onto the eigenvectors of $L$. Analogously, it does always exist a projective measurement that saturates BhB [31]. We note that since we are considering PVM’s the number of outcomes coincides with the dimension of the Hilbert space. Both tasks, measurement, and classical processing, can be treated simultaneously as a hermitian operator $\hat{\Theta}=\sum_{i}\tilde{\Theta}(i)\Pi_{i}$ [23]. The conditions of Eq. (LABEL:condiciones) are now written in terms of the operator $\hat{\Theta}$ as
@@
$\vec{x}=$ $\begin{pmatrix}\hat{\Theta}_{1}\\ \hat{\Theta}_{2}\\ \vdots\\ \hat{\Theta}_{\frac{N(N+1))}{2}}\end{pmatrix},\quad\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix}.$
So we have a system of equations equivalent to the one for a classical probability distribution with $\frac{N(N+1)}{2}$ support points. As before if this system of equations cannot be solved the QBhB variance will go to infinity. On the other hand, if the first $\frac{N(N+1)}{2}-1$ derivatives are linearly independent, for all $n>\frac{N(N+1)}{2}-1$ one has that $n^{th}$order QBhB is equal to $\frac{N(N+1)}{2}-1$ order QBhB. The difference between the quantum and classical cases is that in the former we have the freedom to choose a measurement. This extra freedom implies that the highest relevant QBhB bound –which gives a different bound than lower orders– is lifted in the quantum case because $N-1<\frac{N(N+1)}{2}-1$.
@@
To resolve this apparent contradiction, we explain why the $N$-order quantum Bhattacharyya bound (QBhB) need not be equal to the $(N-1)$-order QBhB. The key point is that, to compute the $N$- and $(N+1)$-order QBhB, different measurements have to be performed, resulting in different probability distributions $P_{N}$ and $P_{N+1}$. This extra freedom in the choice of measurements allows quantum systems to access more non-trivial information. In other words, the maximal order for which the QBhB provides distinct information is higher than in classical scenarios. Thus, since $P_{N}$ and $P_{N+1}$ are different probability distributions, there is no discrepancy with what we presented in Section III.1.
III.3 Illustrative example
In this subsection, we show a case in which we see the lifting in the constraints produced by the quantum nature of the problem, meaning that we see that the quantum bounds are not trivial –giving the same value as bounds from lower orders– for $n>N-1$, being $n$ the order of the BhB employed and $N$ the number of measurement outcomes. As in this example we deal with qubits, $N=2$.
$\rho(0)=\frac{1}{2}\mathbb{I}+\frac{(2\lambda-1)}{2}\sigma_{z},$ (22)
where $\lambda\in(0,1)$ and
where the first element corresponds to the QFI. To finally get the BhB, we take the inverse of the matrix and look at the first element of it. In Fig. 2 we can see the lower bound on the variance of the estimator when using the QCR bound and QBhB. Thus, computing the quantum Bhattacharyya bound gives new non-trivial information even though $n>N-1$.
$\alpha(Q^{-1}\bm{\lambda})\bm{L},$
$P_{\theta}^{\text{CRB}}(\alpha_{2})$ $=1-P_{\theta}^{\text{CRB}}(\alpha_{1})$
where $f=2(\theta-\phi)(\theta+\phi)$ and $\alpha_{i}$ are the eigenvalues of $L_{1}$. The probability distribution that saturates the BhB is
$P_{\theta}^{Bh}(\beta_{2})=$ $1-P_{\theta}^{Bh}(\beta_{1})$
and
Figure 2: QCR and QBhB as a function of $\lambda$ for $\theta=1/2$. One sees that they differ, and as expected QBhB bound will be higher than QCR. In addition, we plot these values of $\lambda$ because the plot is symmetric around $\lambda=0.5$ and also divergent for this same value –since the state will be $\rho\propto\mathbb{I}$ and won’t contain any information on the parameter–.
@@ IV Mach-Zehnder interferometer
Historically, greater attention has been given to the Barankin bounds [27, 28]. In contrast, to the best of our knowledge, Bhattacharyya bounds have not been computed for any physical problem before. In this section, we compute the Bhattacharyya bounds and identify the estimators that achieve these bounds for the paradigmatic Mach-Zehnder interferometer. This example shows that when we move away from the point where the locally unbiased estimators are constructed, the Bhattacharyya bounds provide an improvement in the achievable precision, i.e. the MSE of the estimator that satisfies Eq. (17) is smaller than the MSE of the estimator saturating Cramér-Rao.
@@ IV.0.1 Mach-Zehnder
In the Mach-Zehnder configuration, depicted in Fig. 3, two modes of the electromagnetic field interfere on a balanced beamsplitter, the output beams acquire a relative phase $\theta=\varphi_{a}-\varphi_{b}$, and interfere again on the second beamsplitter. Finally, the photon numbers $n_{a}$ and $n_{b}$ are measured at the output ports.
@@
where $J_{q}(x)$ are the Bessel functions. We compute the CRB, the $2^{\text{nd}}$-order BhB, and the estimators achieving each bound at $\theta_{0}=10^{-3}$. In Fig. 4, we compare the mean squared error (MSE) and bias of these two estimators within an interval of $\theta$ that contains $\theta_{0}$.
One can conclude that the BhB offers a more robust bound; i.e., $\text{MSE}_{\text{Bh}}$ remains flat over an interval of $\theta$, whereas $\text{MSE}_{\text{CR}}$ increases rapidly when $\theta\neq\theta_{0}$. Surprisingly, the behaviour of the bias of both estimators is quite similar. In terms of the bias, there is no clear advantage of using the CRB or Bh2 estimator. Finally, the protocol that saturates the BhB achieves a smaller MSE. Therefore, it may be advisable to use this bound as a figure of merit instead of the CRB in certain specific scenarios.
@@
In Fig. 5, we compute the second-order Bhattacharyya bound and the Cramér-Rao bound as a function of the increasing number of photons. We observe that the Cramér-Rao bound achieves the Heisenberg limit, as expected. Interestingly, the Bhattacharyya bound also reaches the Heisenberg limit, indicating that a tighter bound can still achieve Heisenberg scaling. This result suggests an alternative approach to constructing estimators that can achieve optimal scaling even when the knowledge of the true value of the parameter is sparse. In particular, this demonstrates that it is possible to design estimation protocols that are more robust to uncertainties in prior knowledge while still reaching the Heisenberg limit in precision.
Figure 5: A logarithmic plot of the Cramér-Rao bound and the second-order Bhattacharyya bound is shown. As a reference the black dotted line is proportional to $\frac{1}{n^{2}}$. Both bounds have the same scaling as the dotted line, which means that both bounds attain the Heisenberg limit.